OpenStudy (cutiecomittee123):

solve on the interval of [0,2pi) 2sin^2 3sinx-1=0

OpenStudy (cutiecomittee123):

so far I tried using the quadratic equation but I got numbers that i dont understand how to deal with. Please help!

OpenStudy (anonymous):

You are correct on using the quadratic formula, but before you must do that You must substitute something else for "Sin"

OpenStudy (freckles):

we have a cubic unless there is a type-o in the question

OpenStudy (cutiecomittee123):

well I plugged in the numbers as a=2 b=3 and c=-1

OpenStudy (freckles):

nevermind it isn't a cubic I see \[2\sin^2(3 \sin(x))-1=0\] is that right? I thought I see a cubic because I thought I read: \[2\sin^2(x) 3\sin(x)-1 =0 \text{ which is just } 6 \sin^3(x)-1=0\]

OpenStudy (cutiecomittee123):

okay so that equation can simplify to 6sin^3(x)-1=0??

OpenStudy (freckles):

if you meant: \[2 \sin^2(x)3\sin(x)-1=0?\]

OpenStudy (freckles):

but you wrote \[2 \sin^2(3 \sin(x))-1=0\]

OpenStudy (cutiecomittee123):

yes i was wrong the x was supposed to be included

OpenStudy (cutiecomittee123):

so now what? since its cubic how do i solve it

OpenStudy (freckles):

so the equation really is 2 sin^2(x)3sin(x)-1=0?

OpenStudy (cutiecomittee123):

yes

OpenStudy (freckles):

are you you seem to have thought the question was 2sin^2(x)+3sin(x)-1=0 earlier but I'm not sure what you mean the question to be

OpenStudy (cutiecomittee123):

i am just wondering how to solve this equation

OpenStudy (freckles):

lol I'm very confused what you mean the question to be your solution seems to suggest you wrote the question here wrong

OpenStudy (freckles):

you guys called it a quadratic where you picked a=2 and b=3 and c=-1 this seems to suggest the equation is actually 2sin^2(x)+3sin(x)-1=0 but you are saying the equation is 2sin^2(x)3sin(x)-1=0 just want to be sure

OpenStudy (cutiecomittee123):

The equation is 2sin^2(x)3sin(x)-1 How do I solve this? I tried using the quadratic equation but you added that it is a cubic equation so it cannot e solved using the quadratic equation. Now what? How is this solved then

OpenStudy (freckles):

do you know 2 times 3 is 6? and sin^2(x) times sin(x) is sin^3(x)?

OpenStudy (freckles):

\[2 \sin^2(x) 3 \sin(x)-1=0 \\ 2 \cdot 3 \sin^2(x) \sin(x)-1=0 \\ 6 \sin^3(x)=1 \\ \sin^3(x)=\frac{1}{6}\]

OpenStudy (cutiecomittee123):

Im confused because you posted the same equation multiple times asking me if thats thrright one but you are asking me about two of the same exact equations huh

OpenStudy (freckles):

I really think you meant this equation though: \[2\sin^2(x)+3\sin(x)-1=0\]

OpenStudy (cutiecomittee123):

Okay so that is the eqaution simplified and solved?

OpenStudy (freckles):

what?

OpenStudy (freckles):

2sin^2(x)3sin(x)-1=0 is totally not the same as 2sin^2(x)+3sin(x)-1=0

OpenStudy (freckles):

I want you to tell me the equation I don't want to guess anymore what the equation is

OpenStudy (cutiecomittee123):

Ohh I see! Yeah so there is no addition sign inbetween the equation is 2sin^2(x)3sin(x)-1=0

OpenStudy (cutiecomittee123):

therefore what you did when you got the product sin^3(x)=1/6 is correct I think

OpenStudy (cutiecomittee123):

Sorry for making this confusing

OpenStudy (freckles):

yes if you really meant the equation is 2sin^2(x)3sin(x)-1=0 now take the cube root of both sides

OpenStudy (cutiecomittee123):

okay

OpenStudy (cutiecomittee123):

I get sinx=.55

OpenStudy (cutiecomittee123):

shoot did i just delete those?

OpenStudy (freckles):

\[\sin(x)=\sqrt[3]{\frac{1}{6}} \\ \text{ recall } \sin(x) \text{ has period } 2 \pi \\ \sin(x+2 \pi n)=\sqrt[3]{\frac{1}{6}} \\ \text{ now do } \arcsin( ) \text{ on both sides }\] ok also recall \[\sin(-x+\pi)=\sin(x) \text{ since } \\ \sin(-x) \cos(\pi)+\sin(\pi) \cos(-x) \\ -\sin(x)(-1)+0 \\ \sin(x) \\ \text{ so remember we period } 2\pi \\ \text{ so } \\ \sin(x)=\sin(-x+\pi +2 n \pi )=\sin(-x+\pi(1+2n)) \\ \text{ so you also have to solve } \\ \sin(-x+\pi(1+2n))=\sqrt[3]{\frac{1}{6}}\] but you can do this an s similar way take arcsin( ) of both sides

OpenStudy (freckles):

so basically you have to solve \[\sin(x+2 n \pi)=\sqrt[3]{\frac{1}{6}} \text{ and } \sin(-x+\pi(1+2n))=\sqrt[3]{\frac{1}{6}} \\ \text{ using that \arcsin function } \\ \\ \text{ you have } \\ x+2n \pi=\arcsin(\sqrt[3]{\frac{1}{6}}) \text{ and } -x+\pi(1+2 n)=\arcsin(\sqrt[3]{\frac{1}{6}})\]

OpenStudy (freckles):

now it is just two linear equations to solve

OpenStudy (cutiecomittee123):

so just solve the last two and that will give me the soliutions?

OpenStudy (freckles):

well your interval is [0,2pi) you can pick integer n such that your solutions are in the interval [0,2pi) but yes

OpenStudy (cutiecomittee123):

hmm those are some hard to solve equations how do i even start with x+2npi=arcsin(3sqrt(1/6))

OpenStudy (freckles):

subtract 2n pi on both sides

OpenStudy (cutiecomittee123):

I dont understand how the 2npi part can be manipulated? and huh the arcsin confuses me even more, can i just use wolframalpha or something to solve it?

OpenStudy (cutiecomittee123):

Okay

OpenStudy (cutiecomittee123):

now what??

OpenStudy (cutiecomittee123):

x=arcsin(3sqrt(1/6))-2npi

OpenStudy (freckles):

\[\sin(x+2 n \pi)=\sqrt[3]{\frac{1}{6}} \text{ and } \sin(-x+\pi(1+2n))=\sqrt[3]{\frac{1}{6}} \\ \text{ using that \arcsin function } \\ \\ \text{ you have } \\ x+2n \pi=\arcsin(\sqrt[3]{\frac{1}{6}}) \text{ or } -x+\pi(1+2 n)=\arcsin(\sqrt[3]{\frac{1}{6}}) \\ x=-2 n \pi +\arcsin(\sqrt[3]{\frac{1}{6}}) \text{ or} -x=\arcsin(\sqrt[3]{\frac{1}{6}})-\pi(1+2n) \\ x=-2 n \pi+\arcsin(\sqrt[3]{\frac{1}{6}}) \text{ or } x=-\arcsin(\sqrt[3]{\frac{1}{6}})+\pi(1+2n)\]

OpenStudy (mathmate):

@cutiecomittee123 Could you double check to see if the question should read: 2sin^2\(\color{red}{-}\)3sinx-1=0 or 2sin^2\(\color{red}{+}\)3sinx-1=0

OpenStudy (anonymous):

i am going to go with @mathmate probably a quadratic, and since both of those factor easily

OpenStudy (mathmate):

2sin^2(x)3sin(x)-1=0 does not look like a reasonable expression for a question because it would have simplified to 6sin^3(x)-1=0

OpenStudy (freckles):

err lol I asked her several times @satellite73

OpenStudy (anonymous):

i bet $7 it is one of those two

OpenStudy (cutiecomittee123):

i double checked and they arent subtracted or added they are multiplied together

OpenStudy (anonymous):

then i bet $10 it is a typo

OpenStudy (cutiecomittee123):

Wait no, it is added together wow

OpenStudy (mathmate):

@cutiecomittee123 2sin^2(x)3sin(x)-1=0 does not look like a reasonable expression for a question because it would have simplified to 6sin^3(x)-1=0 Can you post an image of the question, please?

OpenStudy (cutiecomittee123):

Yikes lol maybe i should just reask this question

OpenStudy (anonymous):

yes perhaps using an alias

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

actually what you should do is factor the expression and solve that way